OpenStudy (anonymous):

Enter the Quadratic equation in standard form x+10=3(x-1)^2

jhonyy9 (jhonyy9):

what is (x-1)^2

OpenStudy (anonymous):

x + 10 = 3[x² -2x + 1] x + 10 = 3x² - 6x + 3 0 = 3x² - 6x - x + 3 - 10 3x² - 7x - 7 = 0 <------------standard form